Hello my name is Gabriella,
The problem that I need help figuring out whether it was executed right is
For each equation, find the slope of each equation. If the slope is undefined, state this. The problem is 5x-2=2x-7.
I have tried finding the slope of this equation by adding the like terms 5x+2x=7x and -2+-7=5.
I subtracted 5 from both sides and was left with 2x=0, I then divided 2 on both sides and was left with 0.
Is this equation undefined

Answer by algebrahouse.com(1659)

Hello my name is algebrahouse.com {not really}

5x - 2 = 2x - 7
3x - 2 = -7 {subtracted 2x form both sides}
3x = -5 {added 2 to both sides}
x = -5/3 {divided both sides by 3}
slope = undefined

This is a vertical line.
Whenever you have an equation that starts off "x = " and then a number on the
other side of the equal sign, the graph is a vertical line through that number.

Even though your work was a little off, you were right, when you said the slope
was undefined. The slope of a vertical line is undefined.

For more help from me, visit: www.algebrahouse.com